OpenStudy (anonymous):

Serena saved 20% of the money she made babysitting last week. If she made $65 last week, how much did she save? Show your work.

OpenStudy (chriss):

20% is .2 so you simply multiply 65 time .2 to find out how much she saved.

Parth (parthkohli):

She saved: 20/100 * 65 = 1/5 * 65 = $13 She saved 13 dollars.

OpenStudy (anonymous):

thanks dude!

OpenStudy (anonymous):

or girl..

Parth (parthkohli):

DUDE*

OpenStudy (chriss):

Shouldn't just give the answer.